OpenStudy (anonymous):

what is the absolute value of -8+5? is it 3 or 13 ?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

'cos -8+5 =-3 then the |-3|=3

OpenStudy (anonymous):

|-8+5|=|-3|=3

OpenStudy (anonymous):

how come \[\left| a+b \right|= \left| a \right|+\left| b \right| \] I don't get it

hero (hero):

You must be thinking of \[|a \dot\ b|=|a| \dot\ |b|\]

OpenStudy (anonymous):

so what is \[\left| a+b \right|\] =?

OpenStudy (anonymous):

\[\left| a+b \right| = \left| b+a \right|\] ?

hero (hero):

Yeah, that looks right

OpenStudy (anonymous):

ok thanks!